# Dynamical and structural properties of an absorbing phase transition: a case study from granular systems

### v1

@@ V.1 Derivation of the hydrodynamics of the model
Up to this point, the theory we use is neglecting all spatial dependence. Starting from the Boltzmann equation of our system and the singe-particle microscopic velocity and position distribution, we can derive the corresponding hydrodynamical equations through its velocity moments [150, 151, 152]. Following Refs. 134 and 153, the Boltzmann equation dictating the evolution of the one particle distribution function $f(\bm{r},\bm{v},t)$ in our system is given by:
@@
Molecular chaos was again assumed in Eq. (32) to close the equation for the single velocity distribution. The collisional contribution $J$ takes the usual form of a Master equation [154]. While the first term accounts for collisions increasing the number of particles with velocity $\bm{v}_{i}$, the second term is a loss term for $\bm{v}_{i}$ due to collisions. Both terms are modulated by the relative velocities of the particles along the collision since faster particles collide more often. $\Theta$ is the Heaviside function ensuring that collisions are physical and particles are headed toward each other before colliding. Finally, the factor $1/\alpha^{2}$ comes from the Jacobian associated with the change of variable $\bm{v}^{\prime\prime}\to\bm{v}$ and the change of normal velocity at collision [154].
@@
We also define the number density $n$ and packing fraction field $\phi$; useful quantities related to the (mass) density through $\rho(\bm{r},t)=mn(\bm{r},t)=4m\phi(\bm{r},t)/\pi\sigma^{2}$. By multiplying the Boltzmann equation (Eq. (31)) with different power of $v$ or $u-v$, the evolution equations for the hydrodynamical fields can be obtained. Assuming at first a constant $\Delta$ (no synchronization), we obtain:
@@
where some dependencies on $\bm{r}$ and $t$ have been omitted for ease of reading. Now that the microscopic velocity has been integrated, we will simply denote $\bm{\nabla}$ as the spatial gradient. “::::” is the dyadic dot product understood as: $(\bm{A}:\bm{B})=\sum_{i,j}A_{ij}B_{ij}$. $\bm{\Pi}$ and $\bm{q}$ are respectively the microscopic stress tensors and the microscopic heat flux.
Eqs. (35) are currently undetermined since the currents and $G$ are defined from averages with respect to the microscopic probability distribution. To obtain a useful hydrodynamic description, we must assume or derive a closure in order to relate the currents to the macroscopic fields. One widely used approach to approximate the velocity distribution involves expanding it in terms of the gradient of the hydrodynamical fields, a method commonly known as the Chapman-Enskog expansion [155, 156, 134, 157, 158]. The important assumption being that the spatial and temporal dependence of the distribution of velocity at a given point can be solely described by a functional dependence on the hydrodynamic fields [159]: $f(\bm{r},\bm{v},t)\equiv f[\bm{v}|\rho(\bm{r},t),\bm{u}(\bm{r},t),T(\bm{r},t)]$. The Boltzmann equation is then solved pertubatively order by order, by using the following expansion:
$\begin{split}\bm{\nabla}&\to\varepsilon\bm{\nabla}\\ f&=f^{(0)}+\varepsilon f^{(1)}+\dots\\ \end{split}$ (36)
@@
Performing the Chapman-Enskog expansion at first order leads to the following constitutive relations [134]:
@@
with $\bm{1}$ the unit tensor, $p$ the hydrostatic pressure, $\eta$ and $\zeta$ respectively the shear and bulk viscosity and $\kappa$ and $\mu$ the thermal and diffusive heat conductivity. The latter is 0 at equilibrium and arises in our model from the coupling between density gradient and temperature through the inelastic nature of the collisions [160]. $G\equiv G(\rho(\bm{r},t),T(\bm{r},t))$ has formally the same functional shape as Eq. (12), while $\nu$ is an additional non-equilibrium transport coefficient [123]. Other transport coefficients technically arise for $\tilde{G}$ at the hydrodynamic order $k^{2}$ in the hydrodynamic equations. However, these are omitted here as their contributions are negligible.
Given the typical smallness of the correction due to the presence of a $\Delta$ or $\gamma$ [134, 161] over a wide range of values, we will use the equilibrium hard disk transport coefficients in what follows. We give additional details on the chosen expressions for all the numerical values of these transport coefficients in Appendix E.
While interesting physics is already present in the current state of the theory, synchronization and inactivation of particles are absent from the description and the theory is accurate only deep in the active state. Hence, to properly describe the full dynamics of the system, two modifications would be necessary. First, we would need to introduce the evolution equation for the number of active particles, as given by Eq. (19), including a noise term. Additionally, synchronization effects should be accounted for. However, it would be simpler to assume that synchronization can be incorporated in an ad hoc manner by replacing $\Delta$ in $G$ by the effective $\Delta$ from Eq. (23). After these modifications, the temperature field would be understood as the temperature of the active particles alone. Conceptually, more controlled approximations could be achieved by adding a field to describe the density of synchronized particles. However, we will not pursue this approach. In any case, unless explicitly stated otherwise, the computations will be performed for the $\Delta$ model without synchronization. Since we will mostly place ourselves in the active region at high $\alpha$, we do not expect strong differences between the model with synchronization and the model without synchronization.
@@ V.2 Hydrodynamic matrix and stability analysis
Following Ref. 162, 92, we perform a Fourier transform, with $k$ the momentum variable and define the longitudinal $\delta{u}_{\parallel}(\bm{k},t)=\hat{\bm{k}}\cdot{\delta\bm{u}}(\bm{k},t)$ and transversal velocity field $\delta{u}_{\parallel}(\bm{k},t)=\hat{\bm{k}}_{\perp}\cdot{\delta\bm{u}}(\bm{k},t)$ with $\hat{\bm{k}}$ and $\hat{\bm{k}}_{\perp}$ unit vector in the direction of $\bm{k}$ and perpendicular to $\bm{k}$, respectively. We define $\bm{\Psi}(\bm{k},t)=(\delta\rho,\delta u_{\parallel},\delta u_{\perp},\delta T)$ a vector containing the linearized hydrodynamic field. At first order, Eqs. (35) then read:
@@
The subscript 00 on the pressure and each transport coefficient indicates their evaluation at the homogeneous state. For conciseness, the derivatives are denoted by a subscript: $\partial F/\partial X\equiv F_{X}$. $\eta_{0}^{\parallel}=\eta_{0}+\zeta_{0}$ is the longitudinal viscosity and $\eta_{0}^{\perp}=\eta_{0}$ is the transverse viscosity. As usual, the transverse velocity field decouples from the other modes at linear order. The results obtained are consistent with related ones found in the granular system literature [163, 164, 154, 92] and correctly reduce to equilibrium hydrodynamics in the absence of energy change ($G=0$) and damping [162].
@@
From Eq. (41), we understand that the stability of the homogeneous state is related to the negativity of the real part of the eigenvalues of $\bm{M}$. An analysis of the positivity of the eigenvalues utilizing the Routh–Hurwitz stability criterion [165] demonstrates, as in previous studies of related systems [92], that the active state is stable. Consequently, we conclude that the homogeneous state predicted by the kinetic theory in Sec. IV remains stable against spatial inhomogeneity.
This means that the hydrodynamic theory, which goes beyond the mean field approach, also fails to predict a phase transition since the non-zero homogeneous state is always stable, even at low $\phi$. This justifies a posteriori the introduction of an inactive particle population alongside the temperature field that was done in the previous section in order to achieve a continuous transition. The failure of mean field or hydrodynamic theories without population dynamics is due to the lack of molecular chaos and the breakdown of the equilibrium frequency of collision.
The existence of an inactive population of particles introduces a slow field due to critical slowdown. Therefore, it must be taken into account for the theory to accurately predict the fate of the system. Indeed, as expected in the vicinity of a continuous transition, the mean field long time solution to Eq. (19) is $P_{a}(t)\sim e^{-t/\tau}$ with $\tau\sim(\phi-\phi_{c})^{-1}$, hence the relaxation time scales diverges and the population dynamics must be included in the hydrodynamics theory [166].
@@ V.3 Hydrodynamic modes, dynamic and static correlation functions
While the direct measurement of the evolution of hydrodynamic modes is challenging both experimentally and numerically, spatio-temporal correlations of fluctuations are more accessible [162]. We define the correlation functions we will employ as follows (with a slight abuse of notation):
@@
[…]
@@
The dependence on $\bm{k}^{2}$ of the correlation $\bm{C}$ ensures local conservation of momentum and energy by the noise. At equilibrium, this dependence is necessary because hydrodynamic fields must be locally conserved, implying that noises must be a divergence of a random current. In this non-equilibrium context, where neither momentum nor temperature is conserved, this requirement is less immediate. Nevertheless, the source of stochasticity remains the discrete collisions that the particles making up the fluid undergo, which do conserve momentum. Finally, we note that in the linear regime, the noises are non-multiplicative since their coefficients are evaluated at the homogeneous field values.
@@
It can be integrated to obtain the static correlation function $\bm{S}(\bm{k})$. However, it is usually easier to derive it from the equal-time relation [129]:
@@
with $g$ a complicated function. We recover the equilibrium results in the limit $k\to\infty$, however, at low $k$, we depart from the equilibrium result, notably by having expressions with transport coefficients appearing. Moreover, as seen in Eq. (48), we find strong hyperuniformity [180, 181, 93, 182, 183, 184, 185]: $\bm{S}_{\rho\rho}(k)\sim k^{2}$. This is directly linked to the fact that the noise locally conserves the momentum while the damping does not, and is reminiscent of long-range correlations induced by bulk conservation law and boundary dissipation in self-organized criticality [186, 187, 164, 188, 189, 190, 191, 192, 193, 194, 195]. See Ref. 196 for a recent review. This hyperuniformity in a crystal, allows for the breakdown of the Mermin-Wagner theorem [197, 198, 199, 200].
@@
At equilibrium, when $G=0$, we recover $\bm{S}_{\rho\rho}(k,0)\sim k^{-2}$ as $k\to 0$ [201]. However, in our non-equilibrium setting, this scaling is not verified and at small $k$, $\bm{S}_{\rho\rho}(k,0)$ reaches a plateau. It was postulated in Ref. [93] that this plateau could be used to facilitate the experimental detection of random-organizing hyperuniform fluid. As illustrated by the scaling in Eq. (53), this phenomenon is not exclusive to hyperuniform fluids, but is characteristic of systems with a rapidly relaxing temperature field due to its non conservation. Indeed, it would be observed for a system without damping $\gamma$ (and thus non-hyperuniform) but with a non conserved temperature field ($G_{T}\neq 0$) due to energy change at collision; for example, with $\Delta\neq 0$, $\alpha<1$, and $\gamma=0$.
Note that at equilibrium, due to the fluctuation dissipation theorem, the static correlation functions at small $k$ are easily obtained because $\bm{F}(k,t=0)$ is known. Therefore, the dynamic correlation functions can be found without having to define explicitly a noise [162].
@@ V.4 Numerical results
Figure 10: a) Typical eigenvalues $z$ of the hydrodynamic matrix $\bm{M}(k)$ (Eq. (40)) as a function of $k$. The trivial transverse mode is omitted. Dashed lines are the imaginary part of $z$ while solid ones are their real part. We see three typical regions, I and III where sound cannot propagate ($z$ are purely real) and II, at intermediate $k$, for which sound is present. Inset is a zoom on the region I, normalized by theoretical wavenumber at which we expect a bifurcation $k^{-}$. $\phi=0.3$, $\Delta/\sigma\hat{\tau}=0.015$ and $\gamma\hat{\tau}=0.01$. b) Theoretical (dashed lines) and measured (dots) dynamic structure factor $\bm{S}(k,w)$ as a function of $w$ for different systems. We provide only $w>0$ since the function is even. For each system, $T_{0}/m(\sigma\hat{\tau})^{2}\simeq 0.00025$, $\phi=0.1$ and $\sigma k_{0}=0.046$. “Equilibrium” is a system of hard disks. For each non-equilibrium systems, the temperature is fixed to the same value by varying $\Delta$ and $\gamma$ while keeping $\alpha=0.95$. “$\Delta$ alone” is a system without $\gamma$ ($\Delta/(\sigma\hat{\tau})=0.0008835$ and $\gamma\hat{\tau}=0$). “$\Delta+\gamma$, weak damping” is $\gamma\hat{\tau}=0.0002$ and $\Delta/(\sigma\hat{\tau})=0.001695$. “$\Delta+\gamma$, strong damping” is $\gamma\hat{\tau}=0.007$ and $\Delta/(\sigma\hat{\tau})=0.02$ and the system is starting to enter into the critical region. c) Structure factor for the 4 systems considered in b). Inset is the longitudinal velocity-velocity static correlation function. d) Evolution of the dynamic structure factor as $k$ is varied for the “$\Delta+\gamma$, weak damping” system. d) Evolution of the dynamic structure factor as $k$ is varied for the “$\Delta+\gamma$, strong damping” system. f) Comparison of the theory and the simulations between the model introduced in Ref. 93 and our model for the longitudinal velocity-velocity static correlation function and the transverse one in the inset. In both systems, $\gamma\hat{\tau}=0.006$, $\alpha=1$ and $\phi=0.05$. For our system $\Delta/(\sigma\hat{\tau})=0.046$. For Lei and Ni’s model, $\Delta E/m(\sigma\hat{\tau})^{2}=0.0125$ These choices of parameters lead to approximately the same temperature for both systems $T_{0}/m(\sigma\hat{\tau})^{2}\simeq 0.0158$. In all cases, $N$ varies between $5\times 10^{4}$ and $1.5\times 10^{6}$.
@@
  • “Equilibrium” is a simple Equilibrium hard disk system. Its dynamic structure factor (panel b) exhibits three peaks [202] (only two are visible since we only show positive values of $w$): the Rayleigh peak at $k=0$, related to heat diffusion and to the heat mode, and the Brillouin peaks at $w=\pm w_{s}(k_{0})$, related to propagating sound waves. The structure factor (panel c) and static longitudinal velocity correlation (inset of panel c) are almost flat. The theory works flawlessly in this case [162]. Note that, that at larger $k$, we expect a non-flat structure factor for example, with a shape given by the Ornstein–Zernike theory of liquids [162].
  • $\Delta$ alone” is a non-equilibrium system with $\alpha<1$, $\Delta>0$ but $\gamma=0$ which was thoroughly studied in Ref. 92. Due to the non-conservation of energy in this system at collision (the particles still undergo a ballistic free flight with constant velocity), the Rayleigh peak in the dynamic structure factor begins to disappear. At lower $k$ or lower $\alpha$, it would not be visible at all, and only the Brillouin peak would be observable since sound can still propagate due to the conservation of momentum and density. This shows that even if the temperature field is technically a fast field because it is not conserved, it is still necessary to take it into account in the hydrodynamic description to describe qualitatively the observation since its relaxation time becomes comparable to the ones of the other fields at intermediate $k$. This regime was called quasi-elastic in Ref. 92. Moreover, while at this wavevector $\bm{S}_{\rho\rho}(k,w)$ has an equilibrium like shape with a Rayleigh peak, the corresponding static correlation functions are highly out of equilibrium and display a sharp decrease at low $k$. It should be emphasized that this system is not hyperuniform. Specifically, we can prove that without $\gamma$, contrary to the asymptotics found for $\gamma\neq 0$ in Eqs. (48), the structure factor reaches a finite value. A derivation of this fact is provided in Appendix F.
@@
  • $\Delta+\gamma$, strong damping” is the same system as the previous one, except with significantly increased damping, placing us at the boundaries of the critical region. Here, the sound modes have completely disappeared due to strong non-conservation of momentum. Additionally, the theoretical predictions fail. This failure might be in part attributed to the small population of inactive particles not taken into account in the hydrodynamic slow modes and the slight non-Gaussianity of the velocity distribution. However, a more rigorous analysis indicates that the theoretical framework consistently exhibits deviations with the simulations when $\gamma\neq 0$, although less pronounced, below a specific $k$ where $\gamma$ emerges as the predominant dissipative mechanism. This is why the density structure factor starts to deviate for the “$\Delta+\gamma$, weak damping” model below a given $k$. This is the inverse length at which the effect of $\gamma$ becomes strong compared to the effect of $\alpha$ and $\Delta$. In these cases, the velocity distribution is almost perfectly Gaussian and the discrepancies between simulation and theory cannot be attributed to our Gaussian assumptions. We note again the strong hyperuniformity of the structure factor and of the longitudinal static correlation function. However, the critical hyperuniform scaling $\bm{S}_{\rho\rho}\sim k^{0.5}$ is not observed at all, meaning that we are still far away from the transition.
@@
In the following, we aim to understand the limitations of our model and its region of validity. In Fig. 10d, we present the evolution of the dynamic structure factor for the “$\Delta+\gamma$, weak damping” system as $k$ is varied. At large $k$, the theory works well except at $w=0$, this is likely due to the hydrodynamic approximation requiring a small $k$ since the same departure from the theory is observed for the “equilibrium” system (not shown). Additionally, although not easily visible in the structure factor Fig. 10c, the theory slightly deviates for these wavevectors, resulting in disagreement for the dynamic structure factor. At intermediate $k$, the theory agrees perfectly with the simulations. Decreasing again the wavevector makes us fall into the region dominated by $\gamma$, where neither the structure factor nor the dynamic structure factor is well predicted. We also note that for the smallest $k$, we see the Brillouin’s peak decreasing as we approach region I.

### v2

@@ V Hydrodynamics of the active state
We now derive the hydrodynamic theory of the effective 2D model in its active state. In doing so, we will neglect synchronization effects and any heterogeneity in the population of active particles. This approach allows us to quantitatively predict the structural properties of the active states.
@@ V.1 Derivation of the hydrodynamics of the model
Up to this point, the theory we use is neglecting all spatial dependence. Starting from the Boltzmann equation of our system and the singe-particle microscopic velocity and position distribution, we can derive the corresponding hydrodynamical equations through its velocity moments [pitaevskii2012physical, mazenko2006nonequilibrium, kardar2007statistical]. Following Refs. garzo2018enskog and brey2015hydrodynamics, the Boltzmann equation dictating the evolution of the one particle distribution function $f(\bm{r},\bm{v},t)$ in our system is given by:
@@
Molecular chaos was again assumed in Eq. (32) to close the equation for the single velocity distribution. The collisional contribution $J$ takes the usual form of a Master equation [brilliantov2004kinetic]. While the first term accounts for collisions increasing the number of particles with velocity $\bm{v}_{i}$, the second term is a loss term for $\bm{v}_{i}$ due to collisions. Both terms are modulated by the relative velocities of the particles along the collision since faster particles collide more often. $\Theta$ is the Heaviside function ensuring that collisions are physical and particles are headed toward each other before colliding. Finally, the factor $1/\alpha^{2}$ comes from the Jacobian associated with the change of variable $\bm{v}^{\prime\prime}\to\bm{v}$ and the change of normal velocity at collision [brilliantov2004kinetic].
@@
We also define the number density $n$ and packing fraction field $\phi$; useful quantities related to the (mass) density through $\rho(\bm{r},t)=mn(\bm{r},t)=4m\phi(\bm{r},t)/\pi\sigma^{2}$. By multiplying the Boltzmann equation (Eq. (31)) with different power of $v$ or uvu-v, the evolution equations for the hydrodynamical fields can be obtained. Assuming at first a constant $\Delta$ (no synchronization), we obtain:
@@
[…]
Eqs. (35) are currently undetermined since the currents and $G$ are defined from averages with respect to the microscopic probability distribution. To obtain a useful hydrodynamic description, we must assume or derive a closure in order to relate the currents to the macroscopic fields. One widely used approach to approximate the velocity distribution involves expanding it in terms of the gradient of the hydrodynamical fields, a method commonly known as the Chapman-Enskog expansion [chapman1990mathematical, pinto2025hydrodynamic, garzo2018enskog, chapman1912xii, enskog1917kinetische]. The important assumption being that the spatial and temporal dependence of the distribution of velocity at a given point can be solely described by a functional dependence on the hydrodynamic fields [dorfman2021contemporary]: $f(\bm{r},\bm{v},t)\equiv f[\bm{v}|\rho(\bm{r},t),\bm{u}(\bm{r},t),T(\bm{r},t)]$. The Boltzmann equation is then solved pertubatively order by order, by using the following expansion:
[…]
@@
Performing the Chapman-Enskog expansion at first order leads to the following constitutive relations [garzo2018enskog]:
@@
with $\bm{1}$ the unit tensor, $p$ the hydrostatic pressure, $\eta$ and $\zeta$ respectively the shear and bulk viscosity and $\kappa$ and $\mu$ the thermal and diffusive heat conductivity. The latter is 0 at equilibrium and arises in our model from the coupling between density gradient and temperature through the inelastic nature of the collisions [brey1996model]. $G\equiv G(\rho(\bm{r},t),T(\bm{r},t))$ has formally the same functional shape as Eq. (12), while $\nu$ is an additional non-equilibrium transport coefficient [brey1998hydrodynamics]. Other transport coefficients technically arise for $\tilde{G}$ at the hydrodynamic order $k^{2}$ in the hydrodynamic equations. However, these are omitted here as their contributions are negligible.
Given the typical smallness of the correction due to the presence of a $\Delta$ or $\gamma$ [garzo2018enskog, gomez2022enskog] over a wide range of values, we will use the equilibrium hard disk transport coefficients in what follows. We give additional details on the chosen expressions for all the numerical values of these transport coefficients in Appendix E.
@@ V.2 Hydrodynamic matrix and stability analysis
Following Ref. hansen2013theory, brito2013hydrodynamic, we perform a Fourier transform, with $k$ the momentum variable and define the longitudinal $\delta{u}_{\parallel}(\bm{k},t)=\hat{\bm{k}}\cdot{\delta\bm{u}}(\bm{k},t)$ and transversal velocity field $\delta{u}_{\parallel}(\bm{k},t)=\hat{\bm{k}}_{\perp}\cdot{\delta\bm{u}}(\bm{k},t)$ with $\hat{\bm{k}}$ and $\hat{\bm{k}}_{\perp}$ unit vector in the direction of $\bm{k}$ and perpendicular to $\bm{k}$, respectively. We define $\bm{\Psi}(\bm{k},t)=(\delta\rho,\delta u_{\parallel},\delta u_{\perp},\delta T)$ a vector containing the linearized hydrodynamic field. At first order, Eqs. (35) then read:
@@
The subscript 0 on the pressure and each transport coefficient indicates their evaluation at the homogeneous state. For conciseness, the derivatives are denoted by a subscript: $\partial F/\partial X\equiv F_{X}$. $\eta_{0}^{\parallel}=\eta_{0}+\zeta_{0}$ is the longitudinal viscosity and $\eta_{0}^{\perp}=\eta_{0}$ is the transverse viscosity. As usual, the transverse velocity field decouples from the other modes at linear order. The results obtained are consistent with related ones found in the granular system literature [mcnamara1993hydrodynamic, van1999randomly, brilliantov2004kinetic, brito2013hydrodynamic] and correctly reduce to equilibrium hydrodynamics in the absence of energy change ($G=0$) and damping [hansen2013theory].
@@
From Eq. (41), we understand that the stability of the homogeneous state is related to the negativity of the real part of the eigenvalues of $\bm{M}$. An analysis of the positivity of the eigenvalues utilizing the Routh–Hurwitz stability criterion [nise2020control] demonstrates, as in previous studies of related systems [brito2013hydrodynamic], that the active state is stable. Consequently, we conclude that the homogeneous state predicted by the kinetic theory in Sec. IV remains stable against spatial inhomogeneity.
This means that the hydrodynamic theory, which goes beyond the mean field approach, also fails to predict a phase transition since the non-zero homogeneous state is always stable, even at low $\phi$. This justifies a posteriori the introduction of an inactive particle population in the kinetic theory of Sec. IV in order to achieve a continuous transition.
@@ V.3 Hydrodynamic modes, dynamic and static correlation functions
While the direct measurement of the evolution of hydrodynamic modes is challenging both experimentally and numerically, spatio-temporal correlations of fluctuations are more accessible [hansen2013theory]. We define the correlation functions we will employ as follows (with a slight abuse of notation):
@@
In order to perform the averages $\langle\cdot\rangle$ over different steady state realizations, we need to either introduce a probability distribution for the steady state values of the fields or a dynamical noise to the hydrodynamic equations Eq. (39). We chose the latter approach. When the steady state reached by the hydrodynamical fields is an equilibrium one, the probability distribution of the noises for the linearized equations, assuming Gaussianity, is equivalently found through the Mori-Zwanzig projection method [duran2017general], Einstein’s fluctuation theory or through the fluctuation dissipation theorem [de2006hydrodynamic, vazquez2001fluctuating, landau1980vol]. Out of equilibrium, the choice of an appropriate probability distribution remains ambiguous and, in principle, must be derived from the microscopic dynamics [manacorda2017lattice, manacorda2018lattice, lasanta2015fluctuating, bixon1969boltzmann, brey2009fluctuating, maynar2009fluctuating, bixon1989hard, bouchet2020boltzmann]. However, in our work, we will simply assume that the noise has the same form as in equilibrium, a hypothesis often used with success in granular systems [gradenigo2011fluctuating, pagonabarraga2001randomly]. Therefore, the fluctuating hydrodynamic equations take the form:
@@
The dependence on $\bm{k}^{2}$ of the correlation $\bm{C}$ ensures local conservation of momentum and energy by the noise. At equilibrium, this dependence is necessary because hydrodynamic fields must be locally conserved, implying that noises must be a divergence of a random current. In this non-equilibrium context, where neither momentum nor temperature is conserved, this requirement is less immediate. Nevertheless, the source of stochasticity remains the discrete collisions that the particles making up the fluid undergo, which do conserve momentum. The temperature field, is however not conserved by collisions, and a term in $k^{0}$ is therefore expected. However, such term is usually found to be proportional to the non-gaussianity of the velocity distribution [brey2009fluctuating], which we found to be negligible. Finally, we note that in the linear regime, the noises are non-multiplicative since their coefficients are evaluated at the homogeneous field values.
@@
It can be integrated to obtain the static correlation function $\bm{S}(\bm{k})$. However, it is usually easier to derive it from the equal-time relation [gardiner2009stochastic]:
@@
with $g$ a complicated function. We recover the equilibrium results in the limit $k\to\infty$, however, at low $k$, we depart from the equilibrium result, notably by having expressions with transport coefficients appearing. Moreover, as seen in Eq. (48), we find strong hyperuniformity [lei2024non, kuroda2023microscopic, lei2019hydrodynamics, mukherjee2024anomalous, hazra2024hyperuniformitymasstransportprocesses, PhysRevMaterials.8.104002, anand2025emergent]: $\bm{S}_{\rho\rho}(k)\sim k^{2}$. This is directly linked to the fact that the noise locally conserves the momentum while the damping does not, and is reminiscent of long-range correlations induced by bulk conservation law and boundary dissipation in self-organized criticality [grinstein1990conservation, garrido1990long, van1999randomly, plati2021long, simha2002hydrodynamic, kundu2016long, spohn1983long, dorfman1994generic, grinstein1991generic, bak1992self, giusfredi2024localization]. See Ref. bonachela2009self for a recent review. This hyperuniformity in a crystal, allows for the breakdown of the Mermin-Wagner theorem [maire2024enhancing, kuroda2024long, galliano2023two, keta2024longrangeordertwodimensionalsystems].
@@
At equilibrium, when $G=0$, we recover $\bm{S}_{\rho\rho}(k,0)\sim k^{-2}$ as $k\to 0$ [chaikin1995principles]. However, in our non-equilibrium setting, this scaling is not verified and at small $k$, $\bm{S}_{\rho\rho}(k,0)$ reaches a plateau. It was postulated in Ref. [lei2019hydrodynamics] that this plateau could be used to facilitate the experimental detection of random-organizing hyperuniform fluid. As illustrated by the scaling in Eq. (53), this phenomenon is not exclusive to hyperuniform fluids, but is characteristic of systems with a rapidly relaxing temperature field due to its non conservation. Indeed, it would be observed for a system without damping $\gamma$ (and thus non-hyperuniform) but with a non conserved temperature field ($G_{T}\neq 0$) due to energy change at collision; for example, with $\Delta\neq 0$, $\alpha<1$, and $\gamma=0$.
Note that at equilibrium, due to the fluctuation dissipation theorem, the static correlation functions at small $k$ are easily obtained because $\bm{F}(k,t=0)$ is known. Therefore, the dynamic correlation functions can be found without having to define explicitly a noise [hansen2013theory].
@@ V.4 Numerical results
Figure 10: a) Typical eigenvalues $z$ of the hydrodynamic matrix $\bm{M}(k)$ (Eq. (40)) as a function of $k$. The trivial transverse mode is omitted. Dashed lines are the imaginary part of $z$ while solid ones are their real part. We see three typical regions, I and III where sound cannot propagate ($z$ are purely real) and $I$, at intermediate $k$, for which sound is present. Inset is a zoom on the region I, normalized by theoretical wavenumber at which we expect a bifurcation $k^{-}$. $\phi=0.3$, $(\Delta/\gamma\sigma)=1.5$. b) Theoretical (dashed lines) and measured (dots) dynamic structure factor $\bm{S}(k,w)$ as a function of $w$ for different systems. We provide only $w>0$ since the function is even. For each system the parameters are chosen to keep the kinetic temperature roughly constant, we also take $\phi=0.1$ and $\sigma k_{0}=0.046$. “Equilibrium” is a system of hard disks. For each non-equilibrium systems, the temperature is fixed to the same value by varying $\Delta$ and $\gamma$ while keeping $\alpha=0.95$. $\Delta$ alone” is a system without $\gamma$. $\Delta+\gamma$, weak damping” is $\Delta/(\sigma\gamma)=3.4$. $\Delta+\gamma$, strong damping” is $\Delta/(\sigma\gamma)=2.86$ and the system is starting to enter into the critical region. c) Structure factor for the 4 systems considered in b). Inset is the longitudinal velocity-velocity static correlation function. d) Evolution of the dynamic structure factor as $k$ is varied for the $\Delta+\gamma$, weak damping” system. d) Evolution of the dynamic structure factor as $k$ is varied for the $\Delta+\gamma$, strong damping” system. f) Comparison of the theory and the simulations between the model introduced in Ref. lei2019hydrodynamics and our model for the longitudinal velocity-velocity static correlation function and the transverse one in the inset. In both systems, $\alpha=1$ and $\phi=0.05$. For our system $\Delta/(\sigma\gamma)=7.67$. For Lei and Ni’s model, $\Delta E/m(\sigma\gamma)^{2}=1388$ These choices of parameters lead to approximately the same temperature for both systems. In all cases, NN varies between $5\times 10^{4}$ and $1.5\times 10^{6}$.
@@
  • “Equilibrium” is a simple Equilibrium hard disk system. Its dynamic structure factor (panel b) exhibits three peaks [landau1934structure] (only two are visible since we only show positive values of $w$): the Rayleigh peak at $k=0$, related to heat diffusion and to the heat mode, and the Brillouin peaks at $w=\pm w_{s}(k_{0})$, related to propagating sound waves. The structure factor (panel c) and static longitudinal velocity correlation (inset of panel c) are almost flat. The theory works flawlessly in this case [hansen2013theory]. Note that, that at larger $k$, we expect a non-flat structure factor for example, with a shape given by the Ornstein–Zernike theory of liquids [hansen2013theory].
  • $\Delta$ alone” is a non-equilibrium system with $\alpha<1$, $\Delta>0$ but $\gamma=0$ which was thoroughly studied in Ref. brito2013hydrodynamic. Due to the non-conservation of energy in this system at collision (the particles still undergo a ballistic free flight with constant velocity), the Rayleigh peak in the dynamic structure factor begins to disappear. At lower $k$ or lower $\alpha$, it would not be visible at all, and only the Brillouin peak would be observable since sound can still propagate due to the conservation of momentum and density. This shows that even if the temperature field is technically a fast field because it is not conserved, it is still necessary to take it into account in the hydrodynamic description to describe qualitatively the observation since its relaxation time becomes comparable to the ones of the other fields at intermediate $k$. This regime was called quasi-elastic in Ref. brito2013hydrodynamic. Moreover, while at this wavevector $\bm{S}_{\rho\rho}(k,w)$ has an equilibrium like shape with a Rayleigh peak, the corresponding static correlation functions are highly out of equilibrium and display a sharp decrease at low $k$. It should be emphasized that this system is not hyperuniform. Specifically, we can prove that without $\gamma$, contrary to the asymptotics found for $\gamma\neq 0$ in Eqs. (48), the structure factor reaches a finite value. A derivation of this fact is provided in Appendix F.
@@
  • $\Delta+\gamma$, strong damping” is the same system as the previous one, except with significantly increased damping, placing us at the boundaries of the critical region. Here, the sound modes have completely disappeared due to strong non-conservation of momentum. Additionally, the theoretical predictions fail. This failure might be in part attributed to the small population of inactive particles not taken into account in the hydrodynamic slow modes and the slight non-Gaussianity of the velocity distribution. However, a more rigorous analysis indicates that the theoretical framework consistently exhibits deviations with the simulations when $\gamma\neq 0$, although less pronounced, below a specific $k$ where $\gamma$ emerges as the predominant dissipative mechanism. This is why the density structure factor starts to deviate for the $\Delta+\gamma$, weak damping” model below a given $k$. This is the inverse length at which the effect of $\gamma$ becomes strong compared to the effect of $\alpha$ and $\Delta$. In these cases, the velocity distribution is almost perfectly Gaussian and the discrepancies between simulation and theory cannot be attributed to our Gaussian assumptions. We note again the strong hyperuniformity of the structure factor and of the longitudinal static correlation function. However, the critical hyperuniform scaling $\bm{S}_{\rho\rho}\sim k^{0.5}$ is not observed at all, meaning that we are still far away from the transition.
@@
In the following, we aim to understand the limitations of our model and its region of validity. In Fig. 10d, we present the evolution of the dynamic structure factor for the $\Delta+\gamma$, weak damping” system as $k$ is varied. At large $k$, the theory works well except at $w=0$, this is likely due to the hydrodynamic approximation requiring a small $k$ since the same departure from the theory is observed for the “equilibrium” system (not shown). Additionally, although not easily visible in the structure factor Fig. 10c, the theory slightly deviates for these wavevectors, resulting in disagreement for the dynamic structure factor. At intermediate $k$, the theory agrees perfectly with the simulations. Decreasing again the wavevector makes us fall into the region dominated by $\gamma$, where neither the structure factor nor the dynamic structure factor is well predicted. We also note that for the smallest $k$, we see the Brillouin’s peak decreasing as we approach region I.